# Watermarking Language Models through Language Models

### v1

@@ Abstract
This paper presents a novel framework for watermarking language models through prompts generated by language models. The proposed approach utilizes a multi-model setup, incorporating a Prompting language model to generate watermarking instructions, a Marking language model to embed watermarks within generated content, and a Detecting language model to verify the presence of these watermarks. Experiments are conducted using ChatGPT and Mistral as the Prompting and Marking language models, with detection accuracy evaluated using a pretrained classifier model. Results demonstrate that the proposed framework achieves high classification accuracy across various configurations, with 95% accuracy for ChatGPT, 88.79% for Mistral. These findings validate the and adaptability of the proposed watermarking strategy across different language model architectures. Hence the proposed framework holds promise for applications in content attribution, copyright protection, and model authentication.
Keywords Watermarking  $\cdot$ Large Language Model
@@ 1 Introduction
Language model watermarking refers to the practice of embedding identifiable markers or signatures within the outputs of language models. These watermarks are designed to be imperceptible to end-users but detectable by an authorized system, allowing the origin of the content to be traced back to a particular model or model owner. This process typically involves embedding patterns or codes within the generated text in such a way that they do not affect its quality or meaning while enabling traceability and authentication. Recent advancements in deep learning have opened new opportunities for incorporating these watermarks efficiently within large language models (LLMs).
@@
Recent advancements in watermarking techniques for LLMs have focused on two primary approaches: embedding watermarks directly into generated text and embedding watermarks within the LLM architecture itself. Text-based watermarking methods typically alter token probabilities or manipulate token selections during generation to make watermarks detectable by specific algorithms while maintaining imperceptibility for users. Notable contributions include signal embedding through token manipulation [1, 2], cryptographically inspired undetectable watermarks [3], and adaptive strategies based on token entropy [4]. While these methods achieve various levels of robustness, detectability, and minimal impact on text quality, model-based watermarking techniques embed watermarks within the internal structure of LLMs [5, 6], protecting models against tampering and unauthorized fine-tuning. These methods ensure the integrity of model weights, prevent unauthorized use, and safeguard intellectual property. Despite the progress, challenges remain in achieving dynamic, robust, and imperceptible watermarking that can adapt to different text contexts.
In this work, we propose a novel dynamic watermarking method that leverages the internal capabilities of language models to generate and insert watermarks via prompts during content generation. Unlike existing methods that statically embed watermarks into text or model parameters, our approach dynamically adjusts the watermarking based on the model’s output, ensuring flexibility and robustness across various applications. Our advantages include: (1) We introduce a language model-guided watermarking mechanism that dynamically adapts to the generation process, and (2) We demonstrate that our approach maintains high text quality and detection accuracy without requiring access to the model’s parameters or API.
@@ 2 Related Work
Watermarking in Generated Text. Watermarking in generated text has been an effective approach to identify AI-generated content while protecting intellectual property. Several methods focus on embedding signals directly into the generated outputs of LLMs, making these watermarks detectable by algorithms but imperceptible to humans.
@@
Kirchenbauer et al. [1] introduced a watermarking framework for large language models that embeds signals into generated text, which remain invisible to humans but are algorithmically detectable from a short span of tokens. Their method leverages a randomized selection of "green" tokens to softly promote their use during sampling, thus embedding the watermark with negligible impact on text quality. The watermark detection relies on a statistical test with interpretable p-values, without requiring access to the model’s API or parameters. Zhao et al. [2] proposed the Unigram-Watermark, a method for watermarking LLM-generated text by grouping tokens in a fixed manner. They provide a theoretical framework assessing watermark effectiveness, robustness, and tolerance to text editing and paraphrasing. Experiments demonstrate detection accuracy with small impact on text quality. Wang et al. [8] introduced Codable Text Watermarking for LLMs, enabling watermarks to carry multi-bit customizable information, such as model version or user ID. They use a token-based Balance-Marking method, which employs a proxy language model to split the vocabulary into balanced parts. Their evaluation system includes metrics such as watermark success rate, robustness, and impact on text quality. Christ et al. [3] proposed a cryptographically-inspired method for embedding undetectable watermarks in language model outputs by subtly altering token probabilities. Their approach requires a secret key for detection and ensures no observable change to the output distribution, maintaining the quality of the generated text. The method is based on the existence of one-way functions and is designed to remain undetectable even under adaptive querying. Kuditipudi et al. [9] proposed a watermarking methodology for autoregressive language models that maps tokens to a sequence of random numbers generated with a randomized key. Watermarked text is generated by aligning the output tokens to this random number sequence. The method was applied to multiple models, demonstrating robustness to paraphrasing attacks and reliable detection from a small number of tokens, though detection performance varied across models. Guan et al. [10] introduced CodeIP, a multi-bit watermarking technique for code generated by LLMs that inserts watermarks by manipulating tokens to preserve provenance details such as vendor ID. The approach ensures syntactical correctness by constraining the token prediction process with a type predictor. Experiments across multiple programming languages demonstrated the method’s effectiveness in watermarking while maintaining code correctness. Liu and Bu [4] proposed an adaptive watermarking strategy for LLM-generated text. Their method selectively adds watermarks to high-entropy token distributions, while leaving low-entropy tokens untouched. They adaptively scale the output logits based on semantic embeddings of previously generated text, avoiding the use of fixed token lists. Experiments demonstrate robustness and security, with text quality comparable to un-watermarked models. In contrast to token probability manipulation, Munyer et al. [11] developed DeepTextMark, which embeds watermarks by substituting lexical synonyms within the generated text. This method relies on subtle lexical changes, such as synonym replacements, that preserve the text’s meaning while embedding a watermark, ensuring both robustness and imperceptibility. These diverse approaches illustrate a growing focus on balancing robustness, subtlety, and quality in watermarking AI-generated text.
@@
Watermarking within Language Model Architectures. An emerging area of research focuses on embedding watermarks directly within the parameters of LLMs to safeguard the integrity of the models themselves. Unlike text-based watermarks, these techniques target the internal structure of LLMs, ensuring protection against tampering and misuse at the model level.
@@
Prompt-driven Watermarking. Yao et al. [16] introduced the concept of prompt-driven watermarking for copyright protection. In this approach, a prompt is used to inject watermarks into the output of a language model. The watermark can then be verified by prompting the model again. The technique highlights the flexibility of using prompts to control and verify the presence of watermarks.
Many existing methods for watermarking in generated text and embedding watermarks within language model architectures often require access to the model’s API or internal parameters. This requirement can limit their usability in scenarios with restricted API access or proprietary models. Additionally, these methods tend to be static, meaning they apply the same watermarking strategy regardless of the specific content or context of user requests, which reduces their adaptability. In addition, existing prompt-driven method uses prompts as a mechanism to trigger or verify the watermark for copyright protection and traceability. The concept of prompting a model to watermark itself continuously or systematically across all outputs (without user intervention) is uncharted. To this end, we propose using prompts as a self-guided process initiated by the model.
3 The Proposed Method
In this section, we introduce a novel approach to watermarking language models (LMs) using a combination of language models. The proposed framework includes a prompting language model (Prompting LM), a marking language model (Marking LM), and a detecting language model (Detecting LM). The method is designed to dynamically generate, embed, and detect watermarks within text generated by LM.
Figure 1: The proposed method.
As illustrated in Figure 1, this proposed method leverages multiple LMs to implement a dynamic, adaptive watermarking technique. The prompting LM generates system instructions tailored to the user’s input, the marking LM watermarks the output based on these instructions, and the detecting LM uses binary classification to identify watermarked text.
3.1 Prompting Language Model
The process begins with the input provided by a user, referred to as the [USER REQUEST]. The [USER REQUEST] is used as an input to the Prompting LM, along with a fixed prompt that instructs the model to generate a specific system instruction (denoted as [SYSTEM INSTRUCTION]). The fixed prompt is shown in Figure 2.
Figure 2: The proposed fixed prompt to prompting LM.
An example output [SYSTEM INSTRUCTION] by the prompting LM might specify watermarking strategies such as: "Apply a combination of watermarking strategies that adapt dynamically to the content of the [USER REQUEST]. Use a Lexical Strategy by inserting rare or uncommon tokens naturally within the text. Apply a Semantic Strategy by embedding semantically relevant but less frequent phrases related to deep learning and AI. Additionally, use a Structural Strategy by subtly altering sentence structure to maintain fluency while making the watermark detectable. Ensure the watermarks are evenly distributed throughout the output and can be later identified by a classifier." The instruction may employ different watermarking strategies, such as lexical, semantic, and structural strategies. For instance, the Lexical Strategy may introduce rare or uncommon tokens in the output, the Semantic Strategy may embed semantically relevant but uncommon phrases, and the Structural Strategy may subtly adjust sentence structures in detectable ways.
Mathematically, let the [USER REQUEST] be represented as a sequence of tokens $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$. The Prompting LM generates a [SYSTEM INSTRUCTION] $\mathbf{I}$, which is a sequence of instructions $\mathbf{I}=\{i_{1},i_{2},\dots,i_{m}\}$. The combined input to the next stage is given as $\mathbf{C}=[\mathbf{X},\mathbf{I}]$, where $\mathbf{C}$ is a concatenation of the user request and the system instruction.
3.2 Marking Language Model
The combined input $\mathbf{C}$ is then fed into the Marking LM, which is the language model to be watermarked. This watermarking process is dynamic because the Marking LM applies the watermarking strategy specified in the [SYSTEM INSTRUCTION] in a manner that depends on the content of the [USER REQUEST]. For example, if the instruction specifies the use of rare tokens, these tokens are carefully selected and inserted at contextually appropriate positions within the generated text. This ensures that the tokens blend naturally with the surrounding content, preserving overall fluency and coherence while embedding detectable features. The adaptability of this process allows for diverse watermarking strategies tailored to the varying nature of different requests, ensuring that the watermark remains subtle yet effective.
The watermarked output $\mathbf{Y}=\{y_{1},y_{2},\dots,y_{k}\}$ is a sequence of tokens generated by the Marking LM, where watermarking strategies are applied to modify certain tokens. The dynamic nature of this process ensures that the watermark is content-adaptive rather than static, preventing easily detectable patterns that adversaries might exploit to circumvent watermark detection.
3.3 Detecting Language Model
The Detecting LM is responsible for identifying whether a given text sequence has been watermarked. To achieve this, we utilize a pretrained model with a binary classification head, which is fine-tuned to distinguish between watermarked and non-watermarked text. Let $\mathbf{Y}$ represent the output sequence of tokens from the Marking LM. The pretrained model processes the sequence and outputs a binary prediction $\hat{y}\in\{0,1\}$, where $\hat{y}=1$ indicates a watermarked sequence and $\hat{y}=0$ indicates a non-watermarked sequence. The training objective for the Detecting LM is to minimize the binary cross-entropy loss, which can be expressed as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (1)
where $y_{i}$ is the ground truth label for the $i$-th training sample, indicating whether the sequence is watermarked, and $\hat{y}_{i}$ is the predicted probability that the sequence is watermarked.
4 Experiments
In the pre-print experiments, we tested two main models as the Prompting and Marking LMs: ChatGPT and Mistral. This section reports the data preparation and results for each model configuration.
4.1 Data Preparation
To ensure coverage across a wide range of topics, our data generation process is organized into:[USER REQUEST] generation, [SYSTEM INSTRUCTION] generation, and response generation, followed by labeling for training the detection model.
Category Sample Topics
Learning and Education machine learning, neuroscience, educational games
Science and Technology quantum computing, biotechnology, robotics
Creative Writing and Storytelling fantasy worlds, science fiction, mystery plots
Philosophy and Ethics existentialism, moral relativism, utilitarianism
Health and Wellness mental health awareness, nutrition science, physical fitness
Environment and Nature wildlife conservation, pollution control, climate resilience
History and Culture ancient civilizations, world wars, cultural festivals
Business and Economics supply chain management, stock market, digital currency
Personal Development and Motivation goal setting, time management, self-discipline
Fun and Hypothetical Scenarios alien encounters, time travel possibilities, magic realism

### v2

@@ Abstract
Watermarking the outputs of large language models (LLMs) is critical for provenance tracing, content regulation, and model accountability. Existing approaches often rely on access to model internals or are constrained by static rules and token-level perturbations. Moreover, the idea of steering generative behavior via prompt-based instruction control remains largely underexplored. We introduce a prompt-guided watermarking framework that operates entirely at the input level and requires no access to model parameters or decoding logits. The framework comprises three cooperating components: a Prompting LM that synthesizes watermarking instructions from user prompts, a Marking LM that generates watermarked outputs conditioned on these instructions, and a Detecting LM trained to classify whether a response carries an embedded watermark. This modular design enables dynamic watermarking that adapts to individual prompts while remaining compatible with diverse LLM architectures, including both proprietary and open-weight models. We evaluate the framework over 25 combinations of Prompting and Marking LMs, such as GPT-4o, Mistral, LLaMA3, and DeepSeek. Experimental results show that watermark signals generalize across architectures and remain robust under fine-tuning, model distillation, and prompt-based adversarial attacks, demonstrating the effectiveness and robustness of the proposed approach.
{IEEEImpStatement}
This work introduces a modular, prompt-guided watermarking framework designed to embed and detect imperceptible signals in large language model (LLM) outputs without requiring access to model internals. By leveraging prompt-based instruction generation, the method enables dynamic, context-aware watermarking that adapts to each user input while remaining compatible with black-box and proprietary models. The framework’s flexibility and generality make it applicable to real-world deployment scenarios where control over model weights or decoding processes is not feasible. Extensive evaluation across different Prompting–Marking LM combinations demonstrates that the watermark signal is robust to state-of-the-art attacks. This work advances the state of LLM accountability and verification, providing a practical, scalable, and model-agnostic solution for secure content provenance in generative AI systems.
{IEEEkeywords}
Content authentication, instruction control, large language models, prompt engineering, robust watermarking
@@ 1 Introduction
[…]
Language model watermarking refers to the practice of embedding identifiable markers or signatures within the outputs of language models. These watermarks are designed to be imperceptible to end-users but detectable by an authorized system, allowing the origin of the content to be traced back to a particular model or model owner [1]. This process typically involves embedding patterns or codes within the generated text in such a way that they do not affect its quality or meaning while enabling traceability and authentication. Recent advancements in deep learning have opened new opportunities for incorporating these watermarks efficiently within large language models (LLMs).
Figure 1: Overview of our watermarking scheme. The yellow and blue boxes highlight the watermark generation and watermarking phases respectively.
@@
Recent advancements in watermarking techniques for LLMs have focused on two primary approaches: embedding watermarks directly into generated text and embedding watermarks within the LLM architecture itself. Text-based watermarking methods typically alter token probabilities or manipulate token selections during generation to make watermarks detectable by specific algorithms while maintaining imperceptibility for users. Notable contributions include signal embedding through token manipulation [2, 3], cryptographically inspired undetectable watermarks [4], and adaptive strategies based on token entropy [5]. While these methods achieve various levels of robustness, detectability, and minimal impact on text quality, model-based watermarking techniques embed watermarks within the internal structure of LLMs [6, 7], protecting models against tampering and unauthorized fine-tuning. These methods ensure the integrity of model weights, prevent unauthorized use, and safeguard intellectual property.
Despite the progress, challenges remain in achieving dynamic, robust, and imperceptible watermarking that can adapt to different text contexts. Existing watermarking methods are limited by the reliance on static rules, token-level manipulations, or the need to access model internals such as logits or training weights. Moreover, current approaches often do not take full advantage of the instruction-following capabilities of modern LLMs to adapt watermarking behavior dynamically based on content. These gaps motivate the design of our prompt-guided, multi-LM watermarking framework.
To this end, we propose a novel dynamic watermarking method that leverages the internal capabilities of language models to generate and insert watermarks via prompts during content generation. Unlike existing methods that statically embed watermarks into text or model parameters, the proposed model has the ability to embed watermark signals adaptively per prompt, as guided by an instruction generated in context, rather than relying on a fixed watermarking rule. Thus, our approach dynamically adjusts the watermarking based on user prompts, ensuring flexibility and robustness across various applications.
To operationalize this dynamic approach, we introduce a modular framework that consists of three functionally distinct components: a Prompting LM that generates watermarking instructions from user prompts, a Marking LM that incorporates those instructions to generate watermarked responses, and a Detecting LM trained to classify whether an output carries a watermark. As shown in Figure 1, the process begins when a user submits a prompt to the Marking LM (Step 1). This prompt is forwarded to the Prompting LM, which also receives a fixed prompt template (Step 2). The Prompting LM generates a prompt-specific watermarking instruction based on this template (Step 3). The Marking LM combines the original prompt with the generated instruction to produce a watermarked output (Step 4). Finally, these watermarked outputs, along with their non-watermarked variants are used to train the Detecting LM.
The proposed work makes three main contributions. First, we introduce a prompt-guided watermarking framework that uses a modular, multi-LM design to embed and detect imperceptible signals in generated text, requiring no access to model internals. Second, we demonstrate that the framework is dynamic and adaptive—generating watermarking instructions conditioned on each user prompt—while remaining compatible with diverse, black-box LLMs. Third, we show that the watermarking signals generalize across different model architectures and remain robust under both model-based transformations (fine-tuning and distillation) and prompt-based attacks.
@@ 2.1 Watermarking in Generated Text
Watermarking in generated text has been an effective approach to identify AI-generated content while protecting intellectual property. Several methods focus on embedding signals directly into the generated outputs of LLMs, making these watermarks detectable by algorithms but imperceptible to humans.
@@
Kirchenbauer et al. [2] introduced a watermarking framework for large language models that embeds signals into generated text, which remain invisible to humans but are algorithmically detectable from a short span of tokens. Their method leverages a randomized selection of ”green” tokens to softly promote their use during sampling, thus embedding the watermark with negligible impact on text quality. The watermark detection relies on a statistical test with interpretable p-values, without requiring access to the model’s API or parameters. Zhao et al. [3] proposed the Unigram-Watermark, a method for watermarking LLM-generated text by grouping tokens in a fixed manner. They provide a theoretical framework assessing watermark effectiveness, robustness, and tolerance to text editing and paraphrasing. Experiments demonstrate detection accuracy with small impact on text quality. Wang et al. [9] introduced Codable Text Watermarking for LLMs, enabling watermarks to carry multi-bit customizable information, such as model version or user ID. They use a token-based Balance-Marking method, which employs a proxy language model to split the vocabulary into balanced parts. Their evaluation system includes metrics such as watermark success rate, robustness, and impact on text quality. Christ et al. [4] proposed a cryptographically-inspired method for embedding undetectable watermarks in language model outputs by subtly altering token probabilities. Their approach requires a secret key for detection and ensures no observable change to the output distribution, maintaining the quality of the generated text. The method is based on the existence of one-way functions and is designed to remain undetectable even under adaptive querying. Kuditipudi et al. [10] proposed a watermarking methodology for autoregressive language models that maps tokens to a sequence of random numbers generated with a randomized key. Watermarked text is generated by aligning the output tokens to this random number sequence. The method was applied to multiple models, demonstrating robustness to paraphrasing attacks and reliable detection from a small number of tokens, though detection performance varied across models. Guan et al. [11] introduced CodeIP, a multi-bit watermarking technique for code generated by LLMs that inserts watermarks by manipulating tokens to preserve provenance details such as vendor ID. The approach ensures syntactical correctness by constraining the token prediction process with a type predictor. Experiments across multiple programming languages demonstrated the method’s effectiveness in watermarking while maintaining code correctness. Liu and Bu [5] proposed an adaptive watermarking strategy for LLM-generated text. Their method selectively adds watermarks to high-entropy token distributions, while leaving low-entropy tokens untouched. They adaptively scale the output logits based on semantic embeddings of previously generated text, avoiding the use of fixed token lists. Experiments demonstrate robustness and security, with text quality comparable to un-watermarked models. In contrast to token probability manipulation, Munyer et al. [12] developed DeepTextMark, which embeds watermarks by substituting lexical synonyms within the generated text. This method relies on subtle lexical changes, such as synonym replacements, that preserve the text’s meaning while embedding a watermark, ensuring both robustness and imperceptibility. These diverse approaches illustrate a growing focus on balancing robustness, subtlety, and quality in watermarking AI-generated text.
2.2 Watermarking within Language Model Architectures.
@@
An emerging area of research focuses on embedding watermarks directly within the parameters of LLMs to safeguard the integrity of the models themselves. Unlike text-based watermarks, these techniques target the internal structure of LLMs, ensuring protection against tampering and misuse at the model level.
@@
Yao et al. [17] introduced the concept of prompt-driven watermarking for copyright protection. In this approach, a prompt is used to inject watermarks into the output of a language model. The watermark can then be verified by prompting the model again. The technique highlights the flexibility of using prompts to control and verify the presence of watermarks.
Despite recent progress, challenges remain in achieving dynamic, robust, and imperceptible watermarking that can adapt to different text contexts. Existing methods are often constrained by static rules, token-level perturbations, or the need to access model internals such as logits or training weights. Moreover, prompt-driven approaches have primarily used prompts as triggers or verification signals for watermark detection, not as generative mechanisms for systematic control. The concept of prompting a model to watermark itself continuously and adaptively—without manual intervention—remains largely unexplored. To this end, we introduce a novel framework that treats prompts as self-guided, instruction-bearing controllers. Our approach dynamically generates watermarking instructions via a separate language model and uses these to steer generation in downstream models, even in black-box conditions. This design offers both architectural generality and robustness against attacked conditions such as model fine-tuning and distillation.
3 Proposed Framework and Conceptual Foundations
In this section, we present the proposed framework for prompt-guided watermarking using language models. We first establish a conceptual foundation that formulates watermark embedding as a prompt-induced modulation of language model generation. The theoretical exposition covers the mechanism by which reinforcement-tuned language models respond to instruction prompts, the capacity of such prompts to encode latent watermarking signals, and the robustness of the signal under potential transformations. Following this, we instantiate the theory into a concrete implementation pipeline using a multi-model setup. The remainder of the section proceeds in a top-down manner, beginning with the problem definition and theoretical structure, and transitioning to the model components and system design that realize the watermarking objectives in practice.
3.1 Prompt-Driven Watermarking: Problem Overview
Given a user prompt $X=\{x_{1},x_{2},\dots,x_{n}\}$, our objective is to generate an output sequence $Y=\{y_{1},y_{2},\dots,y_{k}\}$ that (i) is semantically aligned and linguistically fluent, (ii) encodes an imperceptible watermark signal, and (iii) allows reliable detection by a downstream classifier. The framework includes three key components: a Prompting Language Model $\mathcal{P}$ that generates system instructions $I=\{i_{1},i_{2},\dots,i_{m}\}$ based on $X$; a Marking Language Model $\mathcal{M}$ that produces $Y$ conditioned on both $X$ and $I$; and a Detecting LM $f$ trained to classify whether $Y$ contains an embedded watermark or not.
The generation process is defined by the conditional distribution:
$P(Y\mid X,I;\theta)=\prod_{t=1}^{k}P(y_{t}\mid y_{<t},X,I;\theta),$ (1)
where $\theta$ denotes the parameters of $\mathcal{M}$.
Here, the instruction $I$ serves as a soft controller that modulates the token generation process of the Marking LM without altering its internal architecture or parameters. The inclusion of $I$ into the conditioning context $C=[X;I]$ results in a shift in the output distribution $P(Y\mid X,I)$ compared to the unconditional baseline $P(Y\mid X)$. This modulation can be interpreted as an alignment process, where the Prompting LM $\mathcal{P}$ encodes a latent watermarking objective into $I$, and the Marking LM $\mathcal{M}$ interprets and instantiates this objective during generation. Since $\mathcal{M}$ operates autoregressively, the influence of $I$ is propagated token-by-token across the sequence, guiding the trajectory of the output through contextual influence rather than explicit architectural control. This indirect but systematic influence enables the embedding of watermark signals with semantic and structural coherence, while preserving generation fluency. The effectiveness of this alignment hinges on the instruction-following capacity of $\mathcal{M}$, which is typically enhanced through reinforcement learning from human feedback (RLHF), and makes such watermarking strategies feasible even without access to model weights or decoding algorithms.
3.2 Watermark Embedding via Instructional Control
Instruction-tuned language models, particularly those refined through RLHF, exhibit a high degree of generation controllability in response to natural language prompts. This makes them highly amenable to watermark embedding, where instructions are used to induce subtle and consistent patterns into generated text without altering its semantics.
To formally characterize this capacity, we define the expected prompt-induced divergence:
$\mathcal{C}_{\text{instr}}=\mathbb{E}_{X}\left[\mathrm{D_{KL}}\left(P(Y\mid X,I)\parallel P(Y\mid X)\right)\right],$ (2)
which quantifies how much an instruction $I$ can shift the model’s output distribution relative to the unprompted baseline. A larger value of $\mathcal{C}_{\text{instr}}$ indicates higher controllability, and consequently, a greater ability to embed latent signal via instructional prompts.
Although our Prompting LM $\mathcal{P}$ is not explicitly trained to maximize a watermarking objective, its pretrained capacity to synthesize coherent, directive prompts allows it to be used as a surrogate for optimization. We interpret this as:
$I^{*}=\arg\max_{I}\mathbb{E}_{Y\sim P(\cdot\mid X,I)}[\mathcal{U}(Y)],$ (3)
where $\mathcal{U}(Y)$ is a latent utility function encoding the twin objectives of semantic fidelity and watermark alignment. In practice, $\mathcal{P}$ generates $I$ by decoding from a fixed instruction template conditioned on $X$, but the effect behaves as if $I$ were optimized to hint watermark structures to the Marking LM $\mathcal{M}$.
We further interpret the instruction $I$ as a soft injection mechanism into the model’s generative policy. The Marking LM $\mathcal{M}$, being autoregressive, responds to $I$ token-by-token, modulating lexical choice, syntactic form, and semantic framing throughout the generation process. This prompt-induced modulation is non-invasive—it does not involve parameter updates or logit rewrites—but operates entirely through context concatenation. This enables multiple watermarking strategies to coexist: for example, rare word insertion, semantic content reshaping, or stylistic restructuring.
Critically, this design leverages the latent alignment between the Prompting LM $\mathcal{P}$ and the Marking LM $\mathcal{M}$. The system instruction $I$ encodes a latent watermarking policy, which $\mathcal{M}$ realizes through controlled text generation. The resulting watermark is embedded at the distributional level—imperceptible to human readers but consistently detectable under statistical analysis.
Hence, the prompt functions both as a carrier of watermark instructions and as a tool for precise modulation of output features. This duality is central to the feasibility of our approach and highlights the novel use of natural language as a control interface for invisible information embedding in modern language models.
To link this embedding stage to detection, we define the expected watermark signal strength in the generated sequence:
$\mathcal{S}_{\text{wm}}=\mathbb{E}_{Y}\left[\sum_{t=1}^{k}\delta(y_{t})\right],$ (4)
where $\delta(y_{t})$ is a latent scoring function quantifying the alignment of token $y_{t}$ with the watermarking intent specified by $I$. While $\delta$ is not explicitly computed during generation, it is implicitly modeled by the downstream detection classifier. A higher value of $\mathcal{S}_{\text{wm}}$ indicates a stronger embedded signal and increases the likelihood of accurate detection, which we formalize in the next subsection.
Figure 2: System Architecture and Pipeline Realization.
3.3 Watermark Detection as Discriminative Inference
Given that watermark signals are embedded through instruction-induced modulation during generation, we approach detection as a discriminative learning task. Specifically, we train a binary classifier $f(\cdot)$ to distinguish watermarked outputs from non-watermarked text. The classifier is implemented by extending a pretrained language model encoder with a lightweight classification head. Given a sequence $Y=\{y_{1},y_{2},\dots,y_{k}\}$, the model predicts a label $\hat{y}\in\{0,1\}$, where $\hat{y}=1$ denotes a watermarked instance.
The model is trained using binary cross-entropy loss:
$\mathcal{L}_{\text{detect}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (5)
where $y_{i}$ is the ground truth and $\hat{y}_{i}=f(Y_{i})$ is the predicted posterior probability for sample $i$.
This classifier implicitly learns a scoring function $\delta(y_{t})$ that reflects the watermarking signal strength at each token. While this function is never explicitly defined or supervised, the classifier’s attention to distributional features across lexical, semantic, and syntactic dimensions allows it to internalize the statistical regularities induced by the watermark embedding process described earlier.
In summary, watermark detection is treated as a model-driven inference problem, where the classifier approximates a latent watermark pattern through supervised contrast between watermarked and non-watermarked distributions. Its effectiveness depends both on the signal strength embedded during generation and the model’s ability to generalize across text variations while preserving detection fidelity.
3.4 System Architecture and Pipeline Realization
The proposed framework is operationalized as a three-stage pipeline involving a Prompting Language Model $\mathcal{P}$, a Marking Language Model $\mathcal{M}$, and a Detecting Language Model $f$. This modular design supports the end-to-end process of instruction-guided watermark generation, embedding, and detection.
Figure 2 depicts the system-level realization of the framework. Each module corresponds to the roles formalized in Section 3.1: the $\mathcal{P}$ synthesizes a content-aware instruction $I$, the $\mathcal{M}$ generates a watermarked output $Y$ conditioned on the concatenated sequence $[X;I]$, and the $f$ performs post-hoc detection based solely on $Y$. The pipeline begins with a user prompt $X=\{x_{1},x_{2},\dots,x_{n}\}$, which is provided to the $\mathcal{P}$ alongside a fixed system prompt template. This template is designed to elicit a [SYSTEM INSTRUCTION] that specifies watermarking strategies customized to the input. These strategies may include lexical injection (e.g., rare tokens), semantic modulation (e.g., topic-relevant phrasing), or structural variation (e.g., syntax changes). To guide the $\mathcal{P}$ in generating consistent and directive instructions, we append this fixed template to $X$ before decoding. The complete content of the fixed prompt is shown in Figure 3.
Figure 3: The fixed system prompt used for guiding the Prompting LM.
The generated instruction $I$ is treated as a latent controller and is not exposed to the end user. The concatenated sequence $C=[X;I]$ is then provided as input to the Marking LM $\mathcal{M}$, which generates the final output sequence $Y=\{y_{1},y_{2},\dots,y_{k}\}$. The model $\mathcal{M}$ follows its usual autoregressive decoding process, but its behavior is modulated by the embedded instruction $I$. The resulting text $Y$ is intended to remain fluent and coherent while embedding a watermark signature that is statistically detectable but semantically unobtrusive.
To enable post-hoc verification, the Detecting LM $f$ is trained to classify whether a given output $Y$ contains a watermark or not. This is achieved by labeling output sequences as either watermarked or clean and training a binary classifier using a frozen pretrained encoder (e.g., RoBERTa) followed by a lightweight feedforward classification head. The classifier is optimized using the binary cross-entropy loss defined in Eq. 5.
We ensure that training data for $f$ includes both watermarked and non-watermarked instances. Non-watermarked samples are generated by passing only the user prompt $X$ to the Marking LM, while watermarked samples are generated using the full pipeline, including the instruction $I$.
This modular architecture enables flexible substitution of components and supports use cases where the Marking LM is closed-source. Additionally, the use of prompt-based control avoids the need for access to model parameters or inference-time logits, making the approach practical for black-box deployment environments.
4 Experiments and Evaluation
4.1 Dataset Preparation
To evaluate the proposed watermarking framework, we construct a synthetic dataset spanning a broad range of topics and prompt structures. The dataset includes user input prompts, watermarking instructions, and corresponding model-generated responses, both watermarked and non-watermarked.
Following the structure described in Section 3, each sample consists of a user prompt $X$, a system instruction $I$ generated by the Prompting LM $\mathcal{P}$, and a response $Y$ produced by the Marking LM $\mathcal{M}$. The dataset is generated in three stages: (1) user prompt generation, (2) system instruction generation, and (3) response collection with watermark toggling.
Categories and Sample Topics
Learning and Education: machine learning, educational games
Science and Technology: robotics, quantum computing
Creative Writing and Storytelling: fantasy worlds, mystery plots
Philosophy and Ethics: moral relativism, existentialism
Health and Wellness: mental health, physical fitness
Environment and Nature: climate resilience, pollution control
History and Culture: ancient civilizations, world wars
Business and Economics: stock market, digital currency
Personal Development and Motivation: time management, goal setting
Fun and Hypothetical Scenarios: time travel, alien encounters